1.   The number of calls coming into a hotel’s reservation center averages three per minute.

a. Find the probability that no calls will arrive in a given one-minute period.

b. Find the probability that at least two calls will arrive in a given one-minute period.


1
Expert's answer
2022-01-31T13:34:12-0500

Let "X=" the number of calls arriving in a given one minute period: "X\sim Po(\lambda t)"

Given "\lambda=3, t=1"

"\lambda t=3(1)=3"

a.


"P(X=0)=\dfrac{e^{-3}\cdot3^0}{0!}=e^{-3}\approx0.049787"


b.


"P(X\geq 2)=1-P(X=0)-P(X=1)""=1-\dfrac{e^{-3}\cdot3^0}{0!}-\dfrac{e^{-3}\cdot3^1}{1!}=1-4e^{-3}\approx0.800852"